# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ III-A Metrics: Towards Significance Representation
Metrics | Significance Aspect | Properties | Evolution Example | Formula |
AoI | The staleness of source data |
(1) | [36, 37] | |||||||
(2) | [38, 39] | |||||||
@@
We start our review on significance metrics by AoI, and then we introduce its variants, including nonlinear AoI, age of synchronization (AoS), age of incorrect information (AoII), urgency of information (UoI) and the recently proposed goal-oriented tensor (GoT), to describe multi-dimensional significance (e.g., timeliness, status synchronization cost, and environment variation cost). The characteristics of significance metrics reviewed in this survey are listed in Table I. AoI is firstly proposed in [36], where the authors use this metric to find the optimal information transmission rate in vehicular networks. In [37], the authors further use AoI to evaluate the performance of status update systems. Specifically, AoI is defined as the time elapsed since the latest successfully received status update was generated. The instantaneous AoI expression for the system is
@@
where represents the current time and represents the generation time of the currently received information. The evolution of AoI is rather simple, that is, only when a status update is successfully received does AoI update as the age of the newly received update, and otherwise AoI continuously increases in a linear manner. In the SAGSIN, status update messages are exchanged among UAVs, satellites, ships, and ground base stations through wireless channels. High AoI implies long delay and frequent errors of the status update transmission, thereby affecting the significance (or timeliness) of data. Usually, the average AoI or peak AoI is used to measure long-term significance of data. Some research also focuses on the AoI at specific query instants to propose query age of information (QAoI)[44] as an AoI variant. However, the linearly increasing cost of AoI is insufficient in describing the nonlinear significance variation incurred by untimely information. For applications that are more sensitive to time, such as autonomous driving and remote surgery, the performance degradation caused by information aging may not be a linear function of time. For instance, in the state estimation problem of Gaussian linear time-invariant (LTI) systems, if the system is stable, the state estimation error (which directly determine the significance of source data) is a sublinear function of ; if the system is unstable, the state estimation error increases exponentially with [45]. Therefore, it is necessary to evaluate the significance of data which is varying nonlinearly in age. Some scholars have proposed value of information (VoI)[38] (also called nonlinear AoI[39]). Nonlinear AoI is a nonlinear function of AoI, that is
@@
where is a nonlinear penalty function for linear information age. For example, if the penalty function is chosen as exponential function, a possible evolution of nonlinear AoI is shown in Table I. The choice of penalty function is dependent on how the significance varies in data aging. If the significance increases slowly in time, we adopt functions with decreasing derivatives in time, such as logarithmic function; in contrast, if the significance increases rapidly, we adopt functions with increasing derivatives in time, such as exponential function. However, both AoI and nonlinear AoI only reflect the time elapsed from the generation of the current information to its successful reception, without considering whether the receiver correctly estimates the information from the source. That is, they only focus on the timeliness of the currently transmitting information and ignore the real-time status synchronization between the recovered information and the current source status information. This characteristic of AoI (and nonlinear AoI) is also called content-agnostic. However, synchronization of statuses between transceivers also affects the significance of data, since wrong estimation will lead to biased decision on current status and cause severe consequences. For instance, missed detection of the fire will cause considerably large financial loss due to untimely rescue. Therefore, to reflect whether the receiver fully understands the source information, researchers have proposed a series of content-aware significance metrics including AoS, AoII, and UoI. In [40] the authors define a novel metric called AoS, which extends the freshness of information to the synchronization duration of information. Unlike AoI, AoS is calculated as the duration since the latest time of perfect inference of the current process status, that is
@@
where represents the latest time slot that the statuses between transceivers are synchronized. A possible AoS evolution shown in Table I demonstrates that when statuses are synchronized, AoS is equal to zero, and otherwise AoS linearly increases in asynchronized status duration. As a general case of AoS, AoII is defined as a function of AoS. Specifically, in [41], AoII is described as a (non)linear function with regard to time duration multiplying a function reflecting the difference between source status and estimated status, that is
@@
where is the source status at time , is the estimated result of predicted by the receiver, is an increasing time penalty function, and is an information penalty function that reflects the difference between the predicted result of the receiver and the actual status. Since the difference of two identical statuses is zero, the information penalty function should be equal to zero when . The AoII evolution is similar to that of AoS, that is, if the statuses between transceivers are synchronized, AoII is zero; if they are asynchronized, then AoII increases with the time duration of the asynchronized state in a (non)linear manner. Furthermore, [42] proposes a new metric called UoI, which is used to measure the nonlinear time-varying significance of state information by considering the impact of external environment. Specifically, UoI is defined as
@@
where represents the environment states at time . That is, the time penalty function is time-variant according to the current environment state . As shown in Table I, the UoI evolution is rather complex due to time-variant environment. Even the current statuses between transceivers keep asynchronized during time interval , the manners of UoI increase are different in (linear) and (exponential), for the data in the latter time interval are more significant due to more urgent environment. The recent works enrich the concept of significance to propose comprehensive tensor-based metric which unifies the above significance metrics. Specifically, the authors in [43] investigate a novel performance metric called GoT to directly quantify the significance of data considering the impacts of environments, status synchronization, and inherent costs, which is defined as
@@
where is a tensor with three dimensions, namely source status, estimated status, and environment states. Each element of GoT reflects the significance value (i.e., inherent cost) depending on the difference of source and estimated statuses between transceivers and current environment state, as shown in Table I, where the significance values in different time intervals are rather distinct. The authors prove that the GoT metric can reduce to existing metrics such as AoI, AoS, AoII, and UoI under certain conditions, as illustrated in Fig. 5.
@@ III-B Techniques: Towards Significance Protection
This subsection discusses the design of a significance-oriented semantic communication system within SAGSIN. The main focus is on ensuring appropriate sampling and accurate transmission of information with high significance, referred to as “significance protection”. Taking into account the characteristics of SAGSIN, such as large-scale scenarios, highly dynamic channels, and limited device capabilities, we categorize existing research into the design of sampling, coding, and modulation policies, which we term as SPTs. These techniques aim to optimize the significance of transmitted data throughout the signal processing and transmission stages, which are listed in Table II.
@@ III-B1 Significance-oriented sampling policy design
Content-agnostic sampling policy design. Since AoI is a typical timeliness metric, the design of AoI-related sampling policy has been studied widely in the literature. Intuitively, a zero-wait sampling policy, where a new update is sampled just when an acknowledgment (ACK) feedback is received by transmitter, will be naturally AoI-optimal since the newest sampled status update is the freshest and contains the most semantics. However, the authors in [46] point out that zero-wait policy is not always AoI-optimal, especially when the status varies slowly, since most of the status updates are repetitive and thus useless. To solve the AoI-optimal sampling problem for such cases, [46] adopts a penalty function to measure the dissatisfaction level on data staleness, and formulate the age penalty minimization problem as a semi Markov decision process (MDP), which is solved by a divide-and-conquer approach. The zero-wait policy is proven to be optimal when the variance of log-normal distribution for service time is small; in contrast, the policy of waiting is optimal when the variance is large (i.e., a heavy tail distribution) or the penalty function increases rapidly with age growing. Similarly, the AoI-optimal sampling policy under energy constraint is solved in [47]. The choice of nonlinear age penalty and its impact on optimal sampling policy can also be seen in [39]. With regard to AoI-optimal sampling policy for specific distributions of source data, [48] discusses the optimal sampling policies for remote estimation of a Wiener process for different metrics including AoI and MSE. Specifically, the AoI-optimal sampling policy follows a threshold structure, that is, only when AoI reaches a threshold does the sampler generate a new status update for transmission. Meanwhile, the MSE-optimal sampling policy is equivalent to AoI-optimal one only when sampling times are independent with Wiener process; however, when the sampling times are determined by the knowledge of the Wiener process, the AoI-optimal sampling policy will no longer achieve the minimum estimation error as compared to MSE-optimal one. Another similar work is [49], where the optimal sampling policy of a Ornstein-Uhlenbeck source is studied serving as a general case of the study in [48]. The sampling optimization problem is formulated as a time-continuous MDP, and the solution is also proven to be a threshold structure related to instantaneous estimation error (which is related to MSE). The above research focuses on sampling design of single source. For the case of multiple sources, [50] studies the total-peak-AoI-optimal and total-average-AoI-optimal sampling strategies, respectively. To simplify the solution of overall scheduling-sampling policy for multi-source case, the authors firstly prove that a maximum age first (MAF) scheduling policy provides the best age performance, and the optimal sampling policy for each source can be solved easily by a dynamic programming algorithm based on the MAF policy. The total-peak-AoI-optimal sampling policy is proven to be zero-wait, while the total-average-AoI-optimal sampling policy can be solved by water filling, which holds a threshold structure similar to [46, 48, 49]. Content-aware sampling policy design. Although AoI-oriented sampling policy do selects more significant data for transmission, this design is inherently content-agnostic, since the sampler selects only fresh status updates without knowing the content of data. In order to design a content-aware significance-oriented sampling policy, we should resort to AoI variants such as AoS and AoII, etc.
@@
A major content-aware significance metric is AoII, and thus we review AoII-related data sampling optimization. In [40], the AoS-optimal and AoI-optimal sampling rate are respectively solved under a multi-source scenario. As a general case of AoS, the authors in [41] discuss AoII-optimal sampling policy with and without power constraints respectively. Under no power constraint, the optimal policy is “always update”, which also achieves the minimum AoI and MSE simultaneously; however, the optimal policy is complex under a power constraint, which is solved by constraint MDP. [51] further studies AoII-optimal sampling policy under power constraint over an unreliable channel, where a threshold-based policy is proven to be optimal and adopted to reduce the complexity of global optimization problem. The authors in [52] point out the semantic (i.e., significance) characteristics of AoII, and showcase the superiority of significance-oriented AoII-optimal policy as compared to MSE-optimal and AoI-optimal ones. Considering the non-trivial transmission delay which is a fundamental characteristic in the SAGSIN, [53] solves AoII-optimal sampling policy when the status updates experience random delay, which also holds a threshold structure related to the maximum transmission delay. Also, a recent work[54] introduces limited retransmission with resource constraint to enhance the AoII performance.
@@ III-B2 Significance-oriented coding policy design
Considering the large-scale scenarios and highly dynamic channels of the SAGSIN, the implementation of significance-oriented coding policy necessitates the allocation of data rates, striking a trade-off between system efficiency and reliability. In particular, in scenarios involving bad channel conditions or a large volume of significant data, introducing more redundancy can improve the successful transmission probability. However, in contrast, when the channel conditions are favorable or there are less important data, using fewer codewords can still yield satisfactory results. As a widely used retransmission protocol, hybrid automatic repeat request (hybrid ARQ or HARQ), including HARQ with chase combining (HARQ-CC) and HARQ with incremental redundancy (HARQ-IR), accomplishes code rate allocation by dynamically adapting the number of retransmissions or/and the redundancy length in each retransmission, taking into account the factors such as channel conditions, delay constraints, and data significance. The differences among simple ARQ protocol, HARQ-CC protocol, and HARQ-IR protocol are illustrated in Fig. 7.
Coding policy design without considering propagation delay. Most of the works in this field have only considered the transmission delay and neglected the propagation delay, i.e., the only delay elements that affect AoI evolution is the code length. For example, [55] considers a scenario where the sensor transmits the collected data to a central entity over an unreliable link, and coding policies with low latency are designed to protect significant data. HARQ is used for the transmission of updates and the relationship between average timeliness and physical layer decisions (i.e., whether to retransmit the old data or send a new data) is analyzed. The analysis reveals a trade-off between average feedback rate and average timeliness. Specifically, for a constrained set of HARQ code word lengths, refining the code length can improve the average timeliness at the receiver. The work formulates the average age as the objective function and finds out the block allocation vector that minimizes the average age under the constraint of the average feedback rate. The results show that HARQ can greatly outperform ARQ and fixed-length schemes with no retransmission when the size of the incremental redundancy sub-block is properly chosen. The HARQ protocol considered in [55] is called HARQ-IR. Different from the ARQ scheme where the same packet is retransmitted after a failed decoding until ACK reception, the HARQ-IR scheme combines previously received symbols in decoding process to enhance successful decoding probability [70]. In [56], the authors utilize the HARQ-IR protocol in the status update system over a noisy channel. In case of decoding failure at the receiver, IR bits are transmitted to increase the successful decoding probability in the future. If decoding remains unsuccessful for a specified duration, a new status update is transmitted as a replacement. Conversely, when decoding is successful, the transmitter enters an idle state for a certain period after successful transmission, prior to sending a new update. The research focuses on optimizing the code word and IR length for each update, along with the waiting time, with the objective of minimizing the long-term average AoI over the binary symmetric channel (BSC). In recent years, there has been an increase in research focusing on analyzing and comparing AoI based on retransmission protocols for specific state-of-the-art channel coding techniques. The authors in [57] firstly investigate the AoI performance of specific coding schemes. Specifically, the authors analyze the average AoI and energy cost in LDPC coded status update system with and without ARQ using a fixed redundancy scheme. Different coding policies, including the non-ARQ, classical ARQ, truncated ARQ, and truncated HARQ-CC schemes, are analyzed and compared. Among them, truncated ARQ involves repeated transmission of the current update until a maximum number of transmissions or a successful reception is achieved. Truncated HARQ-CC combines failed packets with soft information from the initial transmission for decoding attempts, and thus eliminates the need for additional encoding schemes as it uses the same codeword for retransmission. Additionally, the accumulation effect of signal-to-noise ratio (SNR) in each HARQ retransmission improves the successful decoding probability. These inherent characteristics empower truncated HARQ-CC to strike a balance between AoI and energy consumption, achieving the best average age and moderate energy cost among the considered schemes. As for AoI performance of status update system coded by capacity-achieving codes, the average AoI performance of Polar coded status updates over the additive white Gaussian noise (AWGN) channel is first investigated in [58]. Similar to the conclusion drawn by [57], compared to non-ARQ schemes, the HARQ-CC scheme can effectively reduce the average AoI, especially in low SNR regions. To further optimize the AoI of the system, two methods are proposed to optimize the design SNR and puncture length, both of which are important construction parameters of Polar codes. Coding policy design considering propagation delay. Instead of only concentrating on transmission delay, some research further considers other types of delays (including propagation delay) and specific application scenarios, thereby providing a more precise depiction of the transmission process in the SAGSIN [59, 60, 61, 62]. The authors in [59] conducted a comprehensive analysis of the AoI for two types of HARQ techniques: reactive HARQ and proactive HARQ. [59] derives unified closed-form expressions for the average AoI and average peak AoI for the two techniques under various types of delay, including coding delay, transmission delay, propagation delay, decoding delay, and feedback delay. Based on the derived explicit expressions, an optimization problem is formulated to minimize the AoI by investigating the optimal codeblock assignment strategy in the finite block-length (FBL) regime. The numerical results show that proactive HARQ offers advantages in terms of both age performance and system robustness. In [60], the authors focus on the scenarios characterized by non-trivial propagation delays, such as space communications and satellite communications. The authors study AoI performance in a communication system with non-trivial propagation delays, where status updates are transmitted to the receiver through a binary erasure channel (BEC). In order to mitigate the effect of erasures on timeliness performance, an HARQ-IR scheme with a predetermined maximum number of retransmissions is employed. It is shown that a critical upper threshold related to propagation delay exists, within which retransmissions benefits the AoI. In cases where the propagation delay is longer than the threshold, allocating all the symbols to the first transmission is the optimal coding policy. Following [60], in [61], the author considers a satellite-based IoT system, where IoT devices observe physical processes and transmit status updates to a monitor node through an error-prone channel with significant propagation delay. [61] applies HARQ-IR scheme to the satellite-based IoT system and investigates the age-optimal redundancy allocation problem under reliability constraints. The results also demonstrate the superiority of HARQ-IR in minimizing AoI below a certain propagation delay threshold. In SAGSIN scenarios, [62] presents a novel fast HARQ-IR protocol which omits the decoding and feedback operations of the standard HARQ-IR scheme in the first few rounds. Under the constraint of finite block lengths, it is also demonstrated that the fast HARQ-IR scheme outperforms in the age-optimal rate allocation problem when the propagation delay is shorter than a certain threshold. Taking into account specific channel coding schemes, the authors in [63] first consider a Spinal coded timely status update system based on HARQ with all practical delay elements. An upper bound of average AoI for Spianl codes is derived. Moreover, an HARQ transmission scheme is optimized in two steps to minimize the AoI. Firstly, the authors optimize the puncture pattern of Spinal codes and propose a transmission scheme based on incremental tail transmission puncture (ITTP). Secondly, an optimal transmission scheme is solved based on coarse-grained ITTP, where the number of symbols in each round is refined. Based on the previous work [58], the authors in [64] investigate a Polar coded status update system that accounts for encoding, transmission, propagation, decoding, and feedback delays. The average AoI of the proposed system with different transmission protocols is analyzed. Based on the analysis, the authors focus on the optimization for Polar-coded HARQ design. Specifically, an effective algorithm is devised to optimize the design SNR in code construction and the code length in HARQ-CC transmission. Additionally, a greedy algorithm is utilized to optimize the code lengths for each transmission and the number of transmissions in Polar-coded HARQ-IR.

### v3

@@ III-A Metrics: Towards Significance Representation
Metrics | Significance Aspect | Properties |
Evolution Example | Formula |
AoI |
(1) | [48, 49] | |||||||||
(2) | [50, 51] | |||||||||
@@
We start our review on significance metrics by AoI, and then we introduce its variants, including nonlinear AoI, age of synchronization (AoS), age of incorrect information (AoII), urgency of information (UoI) and the recently proposed goal-oriented tensor (GoT), to describe multi-dimensional significance (e.g., timeliness, status synchronization cost, and environment variation cost). The characteristics of significance metrics reviewed in this survey are listed in Table III.
AoI is firstly proposed in [48], where the authors use this metric to find the optimal information transmission rate in vehicular networks. In [49], the authors further use AoI to evaluate the performance of status update systems. Specifically, AoI is defined as the time elapsed since the latest successfully received status update was generated. The instantaneous AoI expression for the system is
where represents the current time and represents the generation time of the currently received information. The evolution of AoI is rather simple, that is, only when a status update is successfully received does AoI update as the age of the newly received update, and otherwise AoI continuously increases in a linear manner.
In the SAGSIN, status update messages are exchanged among UAVs, satellites, ships, and ground base stations through wireless channels. High AoI implies long delay and frequent errors of the status update transmission, thereby affecting the significance (or timeliness) of data. Usually, the average AoI or peak AoI is used to measure long-term significance of data. Some research also focuses on the AoI at specific query instants to propose query age of information (QAoI)[56] as an AoI variant.
@@
However, the linearly increasing cost of AoI is insufficient in describing the nonlinear significance variation incurred by untimely information. For applications that are more sensitive to time, such as autonomous driving and remote surgery, the performance degradation caused by information aging may not be a linear function of time. For instance, in the state estimation problem of Gaussian linear time-invariant systems, if the system is stable, the state estimation error (which directly determines the significance of source data) is a sublinear function of ; if the system is unstable, the state estimation error increases exponentially with [57]. Therefore, it is necessary to evaluate the significance of data which is varying nonlinearly in age. Some scholars have proposed value of information (VoI)[50] (also called nonlinear AoI[51]). Nonlinear AoI is a nonlinear function of AoI, that is
where is a nonlinear penalty function for linear information age. For example, if the penalty function is chosen as exponential function, a possible evolution of nonlinear AoI is shown in Table III. The choice of penalty function is dependent on how the significance varies in data aging. If the significance increases slowly in time, we adopt functions with decreasing derivatives in time, such as logarithmic function; in contrast, if the significance increases rapidly, we adopt functions with increasing derivatives in time, such as exponential function.
@@
However, both AoI and nonlinear AoI only reflect the time elapsed from the generation of the current information to its successful reception, without considering whether the receiver correctly estimates the information from the source. That is, they only focus on the timeliness of the currently transmitting information and ignore the real-time status synchronization between the recovered information and the current source status information. This characteristic of AoI (and nonlinear AoI) is also called content-agnostic. However, synchronization of statuses between transceivers also affects the significance of data, since wrong estimation will lead to biased decision on current status and cause severe consequences. For instance, missed detection of the fire will cause considerably large financial loss due to untimely rescue. Therefore, to reflect whether the receiver fully understands the source information, researchers have proposed a series of content-aware significance metrics including AoS, AoII, and UoI.
In [52] the authors define a novel metric called AoS, which extends the freshness of information to the synchronization duration of information. Unlike AoI, AoS is calculated as the duration since the latest time of perfect inference of the current process status, that is
@@
where represents the latest time slot that the statuses between transceivers are synchronized. A possible AoS evolution shown in Table III demonstrates that when statuses are synchronized, AoS is equal to zero, and otherwise AoS linearly increases in asynchronized status duration.
As a general case of AoS, AoII is defined as a function of AoS. Specifically, in [53], AoII is described as a (non)linear function with regard to time duration multiplying a function reflecting the difference between source status and estimated status, that is
@@
where is the source status at time , is the estimated result of predicted by the receiver, is an increasing time penalty function, and is an information penalty function that reflects the difference between the predicted result of the receiver and the actual status. Since the difference of two identical statuses is zero, the information penalty function should be equal to zero when . The AoII evolution is similar to that of AoS, that is, if the statuses between transceivers are synchronized, AoII is zero; if they are asynchronized, then AoII increases with the time duration of the asynchronized state in a (non)linear manner.
Furthermore, [54] proposes a new metric called UoI, which is used to measure the nonlinear time-varying significance of state information by considering the impact of external environment. Specifically, UoI is defined as
@@
where represents the environment states at time . That is, the time penalty function is time-variant according to the current environment state . As shown in Table III, the UoI evolution is rather complex due to time-variant environment. Even the current statuses between transceivers keep asynchronized during time interval , the manners of UoI increase are different in (linear) and (exponential), for the data in the latter time interval are more significant due to more urgent environment.
The recent works enrich the concept of significance to propose comprehensive tensor-based metric which unifies the above significance metrics. Specifically, the authors in [55] investigate a novel performance metric called GoT to directly quantify the significance of data considering the impacts of environments, status synchronization, and inherent costs, which is defined as
@@
where is a tensor with three dimensions, namely source status, estimated status, and environment states. Each element of GoT reflects the significance value (i.e., inherent cost) depending on the difference of source and estimated statuses between transceivers and current environment state. As shown in Table III, the significance values in different time intervals are rather distinct. The authors prove that the GoT metric can reduce to existing metrics such as AoI, AoS, AoII, and UoI under certain conditions, as illustrated in Fig. 5.
Remark 1.
The preference of significance metric depends on our demands on system performance. Specifically, content-agnostic metrics can describe the delay aspect (or staleness), while content-aware metrics captures the fidelity aspect (or mismatch). Therefore, we should firstly explore the demand of specific services and secondly choose corresponding metrics. For instance, in a remote control scenario under time-varying environment, UoI may demonstrate the significance of information better than AoII, since remote control services necessitate precise source status information calling for content-aware metrics, and UoI is more suitable for describing time-varying costs. Also, since GoT is a unified significance metric, it is preferable in most services as long as the number of statuses is sufficiently large. In table III, we have briefly compared the application scenarios of each metric.
@@ III-B Techniques: Towards Significance Protection
This subsection discusses the design of a significance-oriented semantic communication system within SAGSIN. The main focus is on ensuring appropriate sampling and accurate transmission of information with high significance, referred to as “significance protection”. Taking into account the characteristics of SAGSIN, such as large-scale scenarios, highly dynamic channels, and limited device capabilities, we categorize existing research into the design of sampling, coding, and modulation policies, which we term as SPTs. These techniques aim to optimize the significance of transmitted data throughout the signal processing and transmission stages, which are listed in Table IV.
@@ III-B1 Significance-oriented sampling policy design
Content-agnostic sampling policy design. Since AoI is a typical timeliness metric, the design of AoI-related sampling policy has been studied widely in the literature. Intuitively, a zero-wait sampling policy, where a new update is sampled just when an acknowledgment (ACK) feedback is received by transmitter, will be naturally AoI-optimal since the newest sampled status update is the freshest and contains the most semantics. However, the authors in [58] point out that zero-wait policy is not always AoI-optimal, especially when the status varies slowly, since most of the status updates are repetitive and thus useless. To solve the AoI-optimal sampling problem for such cases, [58] adopts a penalty function to measure the dissatisfaction level on data staleness, and formulates the age penalty minimization problem as a semi Markov decision process (MDP), which is solved by a divide-and-conquer approach. The zero-wait policy is proven to be optimal when the variance of log-normal distribution for service time is small; in contrast, the policy of waiting is optimal when the variance is large (i.e., a heavy tail distribution) or the penalty function increases rapidly with age growing. Similarly, the AoI-optimal sampling policy under energy constraint is solved in [59]. The choice of nonlinear age penalty and its impact on optimal sampling policy can also be seen in [51].
With regard to AoI-optimal sampling policy for specific distributions of source data, [60] discusses the optimal sampling policies for remote estimation of a Wiener process for different metrics including AoI and MSE. Specifically, the AoI-optimal sampling policy follows a threshold structure, that is, only when AoI reaches a threshold does the sampler generate a new status update for transmission. Meanwhile, the MSE-optimal sampling policy is equivalent to AoI-optimal one only when sampling times are independent with Wiener process; however, when the sampling times are determined by the knowledge of the Wiener process, the AoI-optimal sampling policy will no longer achieve the minimum estimation error as compared to MSE-optimal one. Another similar work is [61], where the optimal sampling policy of a Ornstein-Uhlenbeck source is studied serving as a general case of the study in [60]. The sampling optimization problem is formulated as a time-continuous MDP, and the solution is also proven to be a threshold structure related to instantaneous estimation error (which is related to MSE).
The above research focuses on sampling design of single source. For the case of multiple sources, [62] studies the total-peak-AoI-optimal and total-average-AoI-optimal sampling strategies, respectively. To simplify the solution of overall scheduling-sampling policy for multi-source case, the authors firstly prove that a maximum age first scheduling policy provides the best age performance. Next, the optimal sampling policy for each source can be solved easily by a dynamic programming algorithm based on the maximum age first policy. The total-peak-AoI-optimal sampling policy is proven to be zero-wait, while the total-average-AoI-optimal sampling policy can be solved by water filling, which holds a threshold structure similar to [58, 60, 61].
Content-aware sampling policy design. Although AoI-oriented sampling policy does select more significant data for transmission, this design is inherently content-agnostic, since the sampler selects only fresh status updates without knowing the content of data. In order to design a content-aware significance-oriented sampling policy, we should resort to AoI variants such as AoS and AoII, etc.
@@
A major content-aware significance metric is AoII, and thus we review AoII-related data sampling optimization. In [52], the AoS-optimal and AoI-optimal sampling rates are respectively solved under a multi-source scenario. As a general case of AoS, the authors in [53] discuss AoII-optimal sampling policy with and without power constraints respectively. Under no power constraint, the optimal policy is “always update”, which also achieves the minimum AoI and MSE simultaneously; however, the optimal policy is complex under a power constraint, which is solved by constraint MDP. [63] further studies AoII-optimal sampling policy under power constraint over an unreliable channel, where a threshold-based policy is proven to be optimal and adopted to reduce the complexity of global optimization problem. The authors in [64] point out the semantic (i.e., significance) characteristics of AoII, and showcase the superiority of significance-oriented AoII-optimal policy as compared to MSE-optimal and AoI-optimal ones. Considering the non-trivial transmission delay which is a fundamental characteristic in the SAGSIN, [65] solves AoII-optimal sampling policy when the status updates experience random delay, which also holds a threshold structure related to the maximum transmission delay. Also, a recent work[66] introduces limited retransmission with resource constraint to enhance the AoII performance.
@@ III-B2 Significance-oriented coding policy design
Considering the large-scale scenarios and highly dynamic channels of the SAGSIN, the implementation of significance-oriented coding policy necessitates the allocation of data rates, striking a trade-off between system efficiency and reliability. In particular, in scenarios involving bad channel conditions or a large volume of significant data, introducing more redundancy can improve the successful transmission probability. However, in contrast, when the channel conditions are favorable or there are less important data, using fewer codewords can still yield satisfactory results. A widely used rate allocation method is introducing retransmission protocols, such as automatic repeat request (ARQ) and hybrid ARQ (HARQ) including HARQ with chase combining (HARQ-CC) and HARQ with incremental redundancy (HARQ-IR). ARQ and HARQ protocols accomplish code rate allocation by dynamically adapting the number of retransmissions or/and the redundancy length in each retransmission, taking into account the factors such as channel conditions, delay constraints, and data significance. The differences among simple ARQ protocol, HARQ-CC protocol, and HARQ-IR protocol are illustrated in Fig. 7.
Coding policy design without considering propagation delay. Most of the works in this field have only considered the transmission delay and neglected the propagation delay, i.e., the only delay elements that affect AoI evolution is the code length. For example, [67] considers a scenario where the sensor transmits the collected data to a central entity over an unreliable link, and coding policies with low latency are designed to protect significant data. HARQ is used for the transmission of updates and the relationship between average timeliness and physical layer decisions (i.e., whether to retransmit the old data or send a new data) is analyzed. The analysis reveals a trade-off between average feedback rate and average timeliness. Specifically, for a constrained set of HARQ code word lengths, refining the code length can improve the average timeliness at the receiver. The work formulates the optimization problem with average age as the objective function and finds out the block allocation vector that minimizes the average age under the constraint of the average feedback rate. The results show that HARQ can greatly outperform ARQ protocol and fixed-length coding schemes with no retransmission when the size of the incremental redundancy sub-block is properly chosen. The HARQ protocol considered in [67] is called HARQ-IR (illustrated in Fig. 7(c)). Different from the ARQ protocol where the same packet is retransmitted after a failed decoding until ACK reception, in the HARQ-IR protocol, a different sub-block from the previous packet will be transmitted in each retransmission, and the receiver combines previous sub-blocks with the current sub-block in the decoding process to enhance successful decoding probability[82].
In [68], the authors utilize the HARQ-IR protocol in the status update system over a noisy channel. In case of decoding failure at the receiver, IR bits are transmitted to increase the successful decoding probability in the future. If decoding remains unsuccessful after a specified duration, a new status update is transmitted as a replacement. Conversely, when decoding is successful, the transmitter enters an idle state for a certain period after successful transmission, prior to sending a new update. The research focuses on optimizing the code word and IR length for each update, along with the waiting time, with the objective of minimizing the long-term average AoI over the binary symmetric channel (BSC).
In recent years, there has been an increase in research focusing on analyzing and comparing AoI based on retransmission protocols for specific state-of-the-art channel coding techniques. The authors in [69] firstly investigate the AoI performance of specific coding schemes. Specifically, the authors analyze the average AoI and energy cost in LDPC coded status update system with and without ARQ using a fixed redundancy scheme. Different coding policies, including the non-ARQ, classical ARQ, truncated ARQ, and truncated HARQ-CC protocols, are analyzed and compared. Among them, truncated ARQ (illustrated in Fig. 7(a)) involves repeated transmission of the current update until a maximum number of transmissions or a successful reception is achieved. Moreover, truncated HARQ-CC (illustrated in Fig. 7(b)) combines previous packets with currently received packet for decoding using maximum ratio combining method. In HARQ-CC, additional encoding schemes are unnecessary as the packets are the same for each retransmission. Additionally, the accumulation effect of signal-to-noise ratio (SNR) in each HARQ retransmission improves the successful decoding probability. These inherent characteristics empower truncated HARQ-CC to strike a balance between AoI and energy consumption, achieving the best average age and moderate energy cost among the considered protocols.
As for AoI performance of status update system coded by capacity-achieving codes, the average AoI performance of Polar coded status updates over the additive white Gaussian noise (AWGN) channel is firstly investigated in [70]. Similar to the conclusion drawn by [69], compared to non-ARQ protocols, the HARQ-CC protocol can effectively reduce the average AoI, especially in low SNR regions. To further optimize the AoI of the system, two methods are proposed to optimize the design SNR and puncture length, both of which are important construction parameters of Polar codes.
Coding policy design considering propagation delay. Instead of only concentrating on transmission delay, some research further considers other types of delays (including propagation delay) and specific application scenarios, thereby providing a more precise depiction of the transmission process in the SAGSIN [71, 72, 73, 74]. The authors in [71] conduct a comprehensive analysis of the AoI for two types of HARQ protocols, namely reactive HARQ and proactive HARQ (the latter protocol is illustrated in Fig. 7(d)). Considering various types of delay, including coding delay, transmission delay, propagation delay, decoding delay, and feedback delay, [71] derives unified closed-form expressions for the average AoI and average peak AoI for the two protocols. Based on the derived explicit expressions, an optimization problem is formulated to minimize the AoI by investigating the optimal block assignment strategy in the finite block-length (FBL) regime. The numerical results show that proactive HARQ offers advantages in terms of both age performance and system robustness.
In [72], the authors focus on the scenarios characterized by non-trivial propagation delays, such as space communications and satellite communications. The authors study AoI performance in a communication system with non-trivial propagation delays, where status updates are transmitted to the receiver through a binary erasure channel (BEC). In order to mitigate the effect of erasures on timeliness performance, an HARQ-IR protocol with a predetermined maximum number of retransmissions is employed. It is shown that a critical upper threshold related to propagation delay exists, within which retransmissions benefits the AoI. In cases where the propagation delay is longer than the threshold, allocating all the symbols to the first transmission is the optimal coding policy. Following [72], in [73], the author considers a satellite-based IoT system, where IoT devices observe physical processes and transmit status updates to a monitor node through an error-prone channel with significant propagation delay. [73] applies HARQ-IR protocol to the satellite-based IoT system and investigates the age-optimal redundancy allocation problem under reliability constraints. The results also demonstrate the superiority of HARQ-IR in minimizing AoI below a certain propagation delay threshold. In SAGSIN scenarios, [74] presents a novel fast HARQ-IR protocol which omits the decoding and feedback operations of the standard HARQ-IR in the first few rounds. Under the constraint of finite block lengths, it is also demonstrated that the fast HARQ-IR protocol outperforms in the age-optimal rate allocation problem when the propagation delay is shorter than a certain threshold.
Taking into account specific channel coding schemes, the authors in [75] firstly consider a Spinal coded timely status update system based on HARQ with all practical delay elements. An upper bound of average AoI for Spinal codes is derived. Moreover, the HARQ transmission protocol is optimized in two steps to minimize the AoI. Firstly, the authors optimize the puncture pattern of Spinal codes and propose a transmission scheme based on incremental tail transmission puncturing pattern. Secondly, an optimal HARQ protocol is solved based on coarse-grained incremental tail transmission puncturing pattern, where the number of code symbols (i.e. the sub-block length) in each round is refined.
Based on the previous work [70], the authors in [76] investigate a Polar coded status update system that accounts for encoding, transmission, propagation, decoding, and feedback delays. The average AoI of the proposed system with different transmission protocols is analyzed. Based on the analysis, the authors focus on the optimization for Polar-coded HARQ design. Specifically, an effective algorithm is devised to optimize the design SNR in code construction and the code length in HARQ-CC transmission. Additionally, a greedy algorithm is utilized to optimize the code lengths for each transmission and the number of transmissions in Polar-coded HARQ-IR.